# A Survey on Efficient Inference for Large Language Models

### v2

@@ 6.2.4 Distributed Systems
In order to achieve high throughput, LLM services are commonly deployed on distributed platforms. Recent works have additionally focused on optimizing the performance of such inference services by exploiting distributed characteristics. Notably, observing that the prefilling is compute-intensive and the decoding is memory-intensive, splitwise [261], TetriInfer [262] and DistServe [263] demonstrate the efficiency of disaggregating the prefilling and the decoding steps of a request. In this way, the two distinct stages are processed independently based on their characteristics. SpotServe [264] is designed to provide LLM service on clouds with preemptible GPU instances. SpotServe efficiently handles challenges including dynamic parallel control and instance migration, and also utilizes the auto-regressive nature of LLMs to achieve token-level state recovery. Moreover, Infinite-LLM [265] extends the paged KV cache method in vLLM  [51] to the distributed cloud environment.

### v3

@@ 6.2.4 Distributed Systems
In order to achieve high throughput, LLM services are commonly deployed on distributed platforms. Recent works have additionally focused on optimizing the performance of such inference services by exploiting distributed characteristics. Notably, observing that the computations of prefilling and decoding have interference with each other, splitwise [272], TetriInfer [273] and DistServe [274] demonstrate the efficiency of disaggregating the prefilling and the decoding steps of a request. In this way, the two distinct steps are processed independently based on their characteristics. ExeGPT [287] also adopts such disaggregated architecture, and proposes different strategies with controllable variables to maximize system throughput under certain latency constraints. Llumnix [288] reschedules the requests at runtime for different serving objectives including load balancing, de-fragmentation, and prioritization. SpotServe [275] is designed to provide LLM service on clouds with preemptible GPU instances. SpotServe efficiently handles challenges including dynamic parallel control and instance migration, and also utilizes the auto-regressive nature of LLMs to achieve token-level state recovery. Moreover, Infinite-LLM [276] parallels different parts of the sequence in the attention operator across the data center, to address the challenges when serving extremely long contexts. LoongServe [289] proposes the elastic sequence parallelism to manage the elastic resource demand at the iteration level, reducing the data movement of KV cache via elaborately designed scheduling.